Answer
$0.19$
Work Step by Step
According to the concept of complements, we have:
$P(A)=1-P(\overline{A})$
We apply this concept here:
$P(\text{18-20 years old})=1-P(\text{not 18-20 years old})=1-(0.03+0.23+0.29+0.25+0.01)=1-0.81=0.19$
You need to log in to continue
You can help us out by revising, improving and updating this answer.
Update this answerAfter you claim an answer you’ll have 24 hours to send in a draft. An editor will review the submission and either publish your submission or provide feedback.